OpenStudy (anonymous):

If the angle y is 58 degrees more than angle z, what is the value of angel x?

OpenStudy (anonymous):



OpenStudy (anonymous):

x + z = 180 because it's cyclic quadrilateral y = z + 58

OpenStudy (anonymous):

y = 2z

OpenStudy (anonymous):

2z = z + 58 z = 58 58 + x = 180

OpenStudy (anonymous):

x=122

OpenStudy (anonymous):

thx